Each of 161 newly manufactured items is examined and the number of scratches per item is recorded (the items are supposed to be free of scratches), yielding the following data: Number of scratches per item Observed frequency 12 44 35 33 12 12 8 5 0 123 4 5 6 7 Let X the number of scratches on a randomly chosen item, and assume that X has a Poisson distribution with parameter μ (a) Find an unbiased estimator of μ and compute the estimate for the data. [Hint: E(X) = μ for X Poisson, so E(X) = ?] (Round your answer to two decimal places.) (b) What is the standard deviation (standard error) of your estimator? Compute the estimated standard error. [Hint: σ/ = μ for X Poisson.] (Round your answer to three decimal places.)
Answer #1

Solution-

(a) Unbiased estimator of population mean = sample mean =

= \frac{\sum f* X_{i}}{\sum f}

= 404/161

= 2.5093

(b) Standard deviation of X = 2.50930.5 [ variance = mean = 2.5093 ]

= 1.5841

Standard error of sample mean =

standard deviation / n0.5

= 1.5841/161^0.5

= 0.1248

Answers

TY!